# Smart Predict–then–Optimize Paradigm for Portfolio Optimization in Real Markets

### v1

@@ Abstract
Improvements in return forecast accuracy do not always lead to proportional improvements in portfolio decision quality, especially under realistic trading frictions and constraints. This paper adopts the Smart Predict–then–Optimize (SPO) paradigm for portfolio optimization in real markets, which explicitly aligns the learning objective with downstream portfolio decision quality rather than pointwise prediction accuracy. Within this paradigm, predictive models are trained using an SPO-based surrogate loss that directly reflects the performance of the resulting investment decisions. To preserve interpretability and robustness, we employ linear predictors built on return-based and technical-indicator features and integrate them with portfolio optimization models that incorporate transaction costs, turnover control, and regularization. We evaluate the proposed approach on U.S. ETF data (2015–2025) using a rolling-window backtest with monthly rebalancing. Empirical results show that decision-focused training consistently improves risk-adjusted performance over predict–then–optimize baselines and classical optimization benchmarks, and yields strong robustness during adverse market regimes (e.g., the 2020 COVID-19). These findings highlight the practical value of the Smart Predict–then–Optimize paradigm for portfolio optimization in realistic and non-stationary financial environments.
@@ 1 Introduction
    We systematically evaluate the Smart Predict–then–Optimize paradigm for portfolio optimization under realistic trading constraints and transaction costs, and contrast it with conventional predict–then–optimize approaches.
@@ 2.1 Portfolio Optimization Modeling
In particular, transaction costs and portfolio turnover are commonly modeled through penalty terms in the optimization objective to account for realistic trading frictions Boyd et al. [2017].
@@ 3.1 Smart Predict then Optimize (SPO) based model
Figure 1: Fig1.
@@
Within this unified structure, different portfolio optimization models are instantiated by specifying different objective functions and optimization parameters, collectively denoted by $\Theta$. These instantiations reflect different investment objectives and practical trading considerations. In the following, we consider three representative formulations: MaxReturn, MaxReturn with transaction costs, and MaxReturn with $\ell_{2}$ weight regularization.
@@ 3.1.1 MaxReturnSPO
3.1.2 MaxReturn with transaction costs
To account for practical trading frictions, we extend the MaxReturnSPO framework by explicitly incorporating transaction costs into the portfolio optimization objective. Let $\bm{w}_{t-1}$ denote the portfolio held in the previous period. The decision variable is obtained by solving
@@ 3.1.2 MaxReturn with transaction costs
where $\gamma>0$ controls the transaction costs intensity.
@@ 3.1.3 MaxReturn with $\ell_{2}$ Weight Regularization
where $\hat{\bm{r}}\in\mathbb{R}^{n}$ denotes the predicted return vector, $\bm{w}\in\mathbb{R}^{n}$ the portfolio weight vector, $\mathcal{W}$ the feasible set of portfolios, $\bm{w}_{t-1}$ the portfolio held in the previous period, $\gamma>0$ the transaction costs coefficient, and $\lambda>0$ the regularization parameter controlling the strength of the $\ell_{2}$ penalty.
@@ 3.3.1 Max Sharpe Ratio
As a traditional optimization-based benchmark, we adopt the classical mean–variance portfolio model. Given the realized return vector $\bm{r}$ and covariance matrix $\Sigma$, the oracle portfolio is obtained by solving
$\bm{w}^{\star}=\arg\max_{\bm{w}\in\mathcal{W}}\frac{\bm{r}^{\top}\bm{w}}{\sqrt{\bm{w}^{\top}\Sigma\bm{w}}}$ (15)
@@ 4.2 Backtest Strategy
Figure 3: Fig3.
@@
In all backtesting experiments, a proportional transaction costs of 0.005 is applied to account for realistic trading frictions.
@@ 4.3 Model Configuration
Proportional transaction costs $\gamma$ 0.005
@@
Table 1 and Table 2 summarizes the model configuration used in our experiments. transaction costs and regularization parameters are fixed across all runs, while the robustness radius is varied as $\rho\in\{0.01,0.1\}$. The learning rate and the number of training epochs are selected from predefined ranges using time-series validation, while the batch size and optimizer are kept constant.
@@ 5.1 Overall Backtest Performance
Figure 4: Fig4.
@@
The SPO+ with transaction costs variant also performs competitively, delivering strong risk-adjusted returns while explicitly accounting for trading costs. In contrast, introducing the $\ell_{2}$ weight regularization term leads to a more conservative and diversified allocation, reducing both return and Sharpe ratio while maintaining drawdown levels comparable to other baselines.
@@ 5.2 Performance during the COVID-19 Market Crisis
Overall, decision-focused learning methods exhibit substantial advantages over the PtO baseline during this turbulent period. Both RobustSPO variants and SPO+ with transaction costs achieve significantly higher risk-adjusted performance, as reflected by markedly improved Sharpe and Sortino ratios, while simultaneously reducing maximum drawdowns relative to PtO. In particular, RobustSPO ($\rho=0.1$) and SPO+ with Fee demonstrate strong downside protection, with maximum drawdowns below 10%, compared to over 30% for the PtO baseline.
@@ 5.3 Performance during 2024 Bull Market
Figure 6: Fig6.
@@ 7 Future Work
Several directions remain for future investigation.
First, while the present study adopts proportional transaction costs for tractability and interpretability, incorporating richer market microstructure effects, such as nonlinear market impact, slippage, and no-trade regions, may further improve realism. Investigating how these frictions interact with decision-focused learning objectives remains an important extension.
In addition, richer risk-aware formulations in the optimization layer, including CVaR, downside-risk measures, and drawdown-aware constraints, offer promising avenues to align decision-focused learning with institutional risk management requirements.
Finally, the robust SPO framework considered in this study may be further generalized by adopting adaptive uncertainty sets or distributionally robust Blanchet et al. [2022] decision-focused objectives, enabling systematic handling of regime shifts and prediction uncertainty. Broader empirical validation across asset universes and market conditions would also help assess the generality of the proposed approach.

### v2

@@ Abstract
Improvements in return forecast accuracy do not always lead to proportional improvements in portfolio decision quality, especially under realistic trading frictions and constraints. This paper adopts the Smart Predict–then–Optimize (SPO) paradigm for portfolio optimization in real markets, which explicitly aligns the learning objective with downstream portfolio decision quality rather than pointwise prediction accuracy. Within this paradigm, predictive models are trained using an SPO-based surrogate loss that directly reflects the performance of the resulting investment decisions. To preserve interpretability and robustness, we employ linear predictors built on return-based and technical-indicator features and integrate them with portfolio optimization models that incorporate transaction fee, turnover control, and regularization. We evaluate the proposed approach on U.S. ETF data (2015–2025) using a rolling-window backtest with monthly rebalancing. Empirical results show that decision-focused training consistently improves risk-adjusted performance over predict–then–optimize baselines and classical optimization benchmarks, and yields strong robustness during adverse market regimes (e.g., the 2020 COVID-19). These findings highlight the practical value of the Smart Predict–then–Optimize paradigm for portfolio optimization in realistic and non-stationary financial environments.
@@ 1 Introduction
    We systematically evaluate the Smart Predict–then–Optimize paradigm for portfolio optimization under realistic trading constraints and transaction fee, and contrast it with conventional predict–then–optimize approaches.
@@ 2.1 Portfolio Optimization Modeling
In particular, transaction fee and portfolio turnover are commonly modeled through penalty terms in the optimization objective to account for realistic trading frictions Boyd et al. [2017].
@@ 3.1 Smart Predict then Optimize (SPO) based model
Figure 1: Overview of the Smart Predict–then–Optimize (SPO) training pipeline for portfolio optimization.
@@
Within this unified structure, different portfolio optimization models are instantiated by specifying different objective functions and optimization parameters, collectively denoted by $\Theta$. These instantiations reflect different investment objectives and practical trading considerations. In the following, we consider three representative formulations: MaxReturn, MaxReturn with transaction fee, and MaxReturn with $\ell_{2}$ weight regularization.
@@ 3.1.1 MaxReturnSPO
3.1.2 MaxReturn with transaction fee
To account for practical trading frictions, we extend the MaxReturnSPO framework by explicitly incorporating transaction fee into the portfolio optimization objective. Let $\bm{w}_{t-1}$ denote the portfolio held in the previous period. The decision variable is obtained by solving
@@ 3.1.2 MaxReturn with transaction fee
where $\gamma>0$ controls the transaction fee intensity.
@@ 3.1.3 MaxReturn with $\ell_{2}$ Weight Regularization
where $\hat{\bm{r}}\in\mathbb{R}^{n}$ denotes the predicted return vector, $\bm{w}\in\mathbb{R}^{n}$ the portfolio weight vector, $\mathcal{W}$ the feasible set of portfolios, $\bm{w}_{t-1}$ the portfolio held in the previous period, $\gamma>0$ the transaction fee coefficient, and $\lambda>0$ the regularization parameter controlling the strength of the $\ell_{2}$ penalty.
@@ 3.3.1 Max Sharpe Ratio
As a traditional optimization-based benchmark, we adopt the classical mean–variance portfolio model. Given the historical mean return vector $\bar{\bm{r}}$ and covariance matrix $\Sigma$, the optimal portfolio is obtained by solving
$\bm{w}^{\star}=\arg\max_{\bm{w}\in\mathcal{W}}\frac{\bar{\bm{r}}^{\top}\bm{w}}{\sqrt{\bm{w}^{\top}\Sigma\bm{w}}}.$ (15)
@@ 4.2 Backtest Strategy
Figure 3: Rolling-window backtesting protocol with monthly rebalancing and time-series validation.
@@
In all backtesting experiments, a proportional transaction fee of 0.005 is deducted from portfolio returns based on portfolio turnover, measured by $\|\bm{w}_{t}-\bm{w}_{t-1}\|_{1}$, to account for realistic trading frictions. This transaction fee is applied uniformly to all compared strategies.
@@ 4.3 Model Configuration
Table 1: Overview of compared strategies.
Category Strategies
SoftmaxDFL Softmax-MaxReturn, Softmax-MaxSharpe
Robust SPO RobustSPO ($\rho=0.01$), RobustSPO ($\rho=0.1$)
SPO SPO+, SPO+ with Fee, SPO+ with turnover penalty
PtO baseline PtO Markowitz
Classical baseline MaxSharpe
@@
Proportional transaction fee $\gamma$ 0.005
@@ 5.1 Overall Backtest Performance
Figure 4: Cumulative net asset value (NAV) curves of all compared strategies over the full backtest period (2016–2024).
@@
The SPO+ with fee variant also performs competitively, delivering strong risk-adjusted returns while explicitly accounting for trading costs. In contrast, introducing the $\ell_{2}$ weight regularization term leads to a more conservative and diversified allocation, reducing both return and Sharpe ratio while maintaining drawdown levels comparable to other baselines.
@@ 5.2 Performance during the COVID-19 Market Crisis
Overall, decision-focused learning methods exhibit substantial advantages over the PtO baseline during this turbulent period. Both RobustSPO variants and SPO+ with Fee achieve significantly higher risk-adjusted performance, as reflected by markedly improved Sharpe and Sortino ratios, while simultaneously reducing maximum drawdowns relative to PtO. In particular, RobustSPO ($\rho=0.1$) and SPO+ with Fee demonstrate strong downside protection, with maximum drawdowns below 10%, compared to over 30% for the PtO baseline.
@@ 5.3 Performance during 2024 Bull Market
Figure 6: Cumulative net asset value (NAV) trajectories during the 2024 bull-market period.
@@ 7 Future Work
Several promising directions remain for future research. First, while the proposed framework accounts for decision-aware optimization under predicted returns, incorporating more advanced downside risk control mechanisms—such as distributionally robust optimization and Value-at-Risk (VaR)–type constraints may further enhance robustness against extreme market movements. Second, although the current study adopts simple and stable predictive models, future work may explore the integration of tree-based predictors within the decision-focused learning framework Elmachtoub et al. [2020]. From an empirical finance perspective, tree-based ensemble models have been shown to exhibit strong robustness in noisy return prediction tasks. For example, Krauss et al. Krauss et al. [2017] demonstrate that random forests and gradient-boosted trees achieve competitive and often superior out-of-sample performance compared to deep neural networks in statistical arbitrage settings, where financial signals are characterized by low signal-to-noise ratios. This robustness makes tree-based predictors a promising direction for future integration into decision-focused portfolio optimization frameworks. Finally, despite the inclusion of diversification-promoting mechanisms, achieving an optimal balance between portfolio diversification and return maximization remains an open challenge. Developing principled approaches to jointly model this trade-off within decision-focused portfolio optimization frameworks constitutes an important direction for future research.